# You May Use the Same Channel Knowledge Map for Environment-Aware NLoS Sensing and Communication

### 17 Mar 2026 v3

@@ III Environment-aware ISAC via CKM
where $\mathcal{M}$ represents the CKM, mapping a two-dimensional UE location vector to a channel vector of dimension $D$. CKM exploits these historical pairs $(\mathbf{Q},\mathbf{W})$ to infer new channel knowledge without requiring an explicit characterization of the mapping function $f(\cdot)$ nor the environment $E$. Without loss of generality, CKM captures the statistical relationship between UE locations $\mathbf{q}$ and channel knowledge $\mathbf{w}$, represented as a conditional probability distribution $P_{\mathcal{M}}(\mathbf{w}|\mathbf{q})$. This relationship is learned from the historical data $(\mathbf{Q},\mathbf{W})$ through various methods, such as interpolation, machine learning, or other data-driven techniques, depending on the specific implementation of CKM.
@@
where $h(\cdot)$ is the function taking the composition of functions $f(\cdot)$ and $g(\cdot)$. From (22), the historic data $(\mathbf{Q},\mathbf{W})$ can effectively replace the requirement to explicitly reconstruct the environment $E$.
CKM is obtained by learning from the historical data $(\mathbf{Q},\mathbf{W})$, represented by the mapping $\mathcal{M}$ as shown in (19). Different from explicitly reconstructing the physical environment $E$ from the historic data $(\mathbf{Q},\mathbf{W})$, which is usually computationally prohibitive and unnecessary, CKM serves as an implicit, data-driven representation of the environment $E$. It directly learns the functional “location-to-channel” mapping from historical data, effectively capturing the necessary environment prior knowledge without the overhead of an explicit physical map reconstruction. Therefore, the observation model (22) can be further written as
@@ IV-A Relationship between sensing response and communication channel
A target that reflects signals to all directions is analogous to a virtual communication UE at the same location. This analogy arises from the physical similarity in signal interaction: the target receives the transmitted signal from the BS, attenuates it with a reflection coefficient $\beta$, and reflects it back to the BS. This process mirrors the behavior of a virtual UE that receives the downlink signal and retransmits it via an uplink channel, as illustrated in Fig. 4.
@@ IV-B CADM-enabled sensing
To model the dependence of these parameters on $\mathbf{q}$, CADM employs an FCNN to learn the mapping from the UE location $\mathbf{q}$ to the mean and variance parameters in the Gaussian model in (36). For example, if a two-layer FCNN is used with the input being the UE location $\mathbf{q}\in\mathbb{R}^{2}$. The forward process is defined as
@@
Then, the log-likelihood of the sensing channel knowledge can be written as (40) shown on the top of the next page, where the second equation is due to (32). In (38) and (39), it assumes that for a given path $l^{\prime}$, the angle $\theta_{l^{\prime}}$ and delay $\tau_{l^{\prime}}$ are statistically independent. This assumption is based on the underlying physics of multipath propagation in complex environments. From a physical standpoint, the parameters are determined by different geometric properties: The AoA is primarily dictated by the geometric orientation of the final scatterer relative to the receiver array. The path delay, in contrast, is a function of the entire path length traveled by the signal, which may involve multiple different scattering objects before reaching that final scatterer. In the NLoS multipath environment, the factors contributing to the total path length (and thus delay) are numerous and distinct from the geometric angle of the final hop. Therefore, in complex environments, the statistical relationship between the total path length and the final arrival angle is generally weak. It is also noted that the CKM framework is not fundamentally restricted by this assumption. The model could be extended to a full multivariate Gaussian distribution.
@@
where $\mathbf{a}\in\mathbb{R}^{4L^{\prime}\times 1}$ is a selection vector with a $1$ at the position corresponding to $\mu_{\theta,l_{T}}(\mathbf{x})$ and 0 elsewhere, $\mathbf{W}_{1}$, $\mathbf{W}_{2}$, $\mathbf{b}_{1}$ are FCNN weights and biases, and $\sigma^{\prime}(\cdot)$ is the derivative of the activation function (such as ReLU). These closed-form expressions enable computation of $\nabla_{\mathbf{x}}(-\ln P(\hat{\mathbf{z}}|\mathbf{x}))$, facilitating optimization of $\mathbf{x}$ via (12). Alternatively, these gradients can be computed efficiently using automatic differentiation in PyTorch by enabling gradient tracking for the location input, calling $backward()$ on the output, and then accessing the input’s .grad.grad attribute. Besides, the $torch.autograd.grad()$ function provides a more direct alternative for this computation.
As a summary, to estimate the location $\mathbf{x}$ of the sensing target for the challenging NLoS environment, parameters such as AoD, AoA, and delay for each path are firstly estimated based on the echo signal $\mathbf{y}(t)$ in (6) using well-established algorithms (e.g., MUSIC, ESPRIT, or compressed sensing [3]), forming the observation vector $\hat{\mathbf{z}}$. The methods cited, such as MUSIC or ESPRIT, are mentioned only as well-established, high-resolution examples, not as mandatory components. For dynamic scenarios involving high-speed targets, where low latency and real-time capability are required, our framework can be seamlessly integrated with various estimation algorithms (e.g., FFT-based) specifically designed for high-mobility tracking. The core contribution of this work lies in intelligently utilizing the outputs $\hat{\mathbf{z}}$ of any such algorithms for NLoS localization. Then, substituting $\hat{\mathbf{z}}$ into (41) allows computation of the gradient, enabling estimation of the target location $\mathbf{x}$ via gradient descent in (12). Note that while the algorithm is presented for the most challenging NLoS scenarios, it is applicable for the LoS case as well. The pseudo-code of CKM-enabled environment-aware sensing is elaborated in Algorithm 1.
@@ V Simulation Results
The simulation is conducted in a 100 m $\times$ 100 m square area, where a BS is located at $(50,0)m$. Besides, there are forty scatterers randomly distributed within the area, and one target is placed at a random location for each sensing process. To model the statistical nature of a dynamic environment, we incorporated small random location variations for the scatterers during the data generation process. This ensures that the training data reflects statistical fluctuations, allowing the FCNN to learn a meaningful location-dependent mean and variance for the channel parameters, rather than just a fixed value. The locations of the BS, scatterers, and target are illustrated in Fig. 5, where the BS is marked with a red triangle, scatterers with green circles, and the sensing target with a blue circle. Two sensing scenarios are considered:
@@
To construct the CKM, we use an FCNN as shown in Fig. 6. It consists of five layers: an input layer with 2 neurons (corresponding to the 2D coordinates of UE location $\mathbf{q}$), three hidden layers with $128\times 256\times 128$ neurons using LeakyReLU activation with a negative slope of 0.01, and an output layer with $4L^{\prime}$ neurons, where $L^{\prime}=5$ is the number of Gaussian components in (36). The output includes the angle mean $\mu_{\theta,l^{\prime}}$, the angle variance $\sigma_{\theta,l^{\prime}}^{2}$, the delay mean $\mu_{\tau,l^{\prime}}$, and the delay variance $\sigma_{\tau,l^{\prime}}^{2}$ for each path $l^{\prime}$. Besides, the training dataset is generated by sampling 10,000 points uniformly within the 100 m $\times$ 100 m area. This dataset was empirically found to be sufficient for the FCNN to converge and was partitioned into a training set (80%) and a validation set (20%). Note that in practical deployments, it is unnecessary to physically collect all 10,000 samples. As mentioned in Section I, advanced CKM construction methods based on generative AI [8, 12] can complete a large-scale dataset from only a small fraction of real-world measurements, significantly reducing the data collection burden. For each sampled UE location in the training dataset, the groundtruth channel is generated by considering all 40 possible single-bounce paths via the scatterers. The delay for each path is calculated based on its total geometric length, which is the distance from the BS to the scatterer and then to the UE. The complex gain of each path is determined by two main factors: the path-length-dependent attenuation and a unique random reflection coefficient assigned to each scatterer. Although 40 paths are generated, the CADM is designed to learn the $L^{\prime}=5$ dominant ones. Therefore, a path selection mechanism is applied: the 5 paths with the highest power, as determined by their path gains, are chosen. The angles and delays of these 5 selected paths constitute the groundtruth data used for training the FCNN and the training loss is computed as the mean squared error (MSE) between the predicted and true channel parameters. The network was trained for 500 epochs using the Adam optimizer, with an initial learning rate of 0.01 and a batch size of 128.
@@
Fig. 7 presents the RMSE for the location estimation of the sensing target with different methods vs. angle estimation error $\sigma_{\theta}$ and $\sigma_{\phi}$, with fixed delay std $\sigma_{\tau}=20$ ns. It can be seen that the CKM-based methods consistently demonstrate significantly lower RMSE than geometry-based approaches across all angle error levels, by orders of magnitude. Such dramatic performance gains are attributed to the utilization of CADM, which exploits a rich set of NLoS path information, including AoA, AoD and delay. In contrast, even in the mixed LoS/NLoS scenarios, the Geometry-based method relies solely on the LoS path’s angle and delay, where angle errors severely degrade performance due to their tight coupling with location as (2). Even with accurate delay estimates, RMSE increases sharply as angle error grows. For the challenging NLoS scenario, the geometry-based method fails to give any meaningful estimation for the sensing target location, since it erroneously assumes the shortest-delay NLoS path as LoS, leading to substantial localization errors. CKM-based methods, by leveraging multiple paths, decouple angle and delay contributions, resulting in much slower RMSE degradation with increasing angle error, as precise delay measurements mitigate the impact of angle inaccuracies. In the pure NLoS scenario, the CKM-based method, employing the proposed generalized sensing model as (6) with up to $L^{\prime 2}$ non-reciprocal paths, outperforms CKM-based NLoS_conv, which uses a conventional model as (5) with only $L^{\prime}$ reciprocal paths. The reduced path diversity in the conventional model limits its information, yielding higher RMSE. Surprisingly, CKM-based NLoS achieves even lower RMSE than CKM-based LoS, especially when the angle errors are large. This seems to be counterintuitive at first glance, but is expected because NLoS paths, with distinct AoA, AoD and delay, provide three parameters per path, compared to two for LoS paths (AoA = AoD, delay). The increased path count ($L^{\prime 2}$ vs. $L^{\prime}$) and richer information per path in NLoS scenarios enhance, when properly exploited as by CKM, robustness to angle errors, enabling superior localization accuracy despite the absence of a LoS path. Similar observations can be obtained from Fig. 8, which plots the sensing RMSE as a function of the delay estimation error $\sigma_{\tau}$.
@@
Following the MSE analysis, the theoretical performance bounds are explored through the CRLB to highlight localization performance limits. CRLB is analyzed vs. angle error (with delay error std $\sigma_{\tau}=20$ ns) and delay error $\sigma_{\tau}$ (with angle error std $\sigma_{\theta}=\sigma_{\phi}=0.10$ rad) as shown in Fig. 9 and 10, respectively. Geometry-based NLoS is excluded, as unknown scatterer location prevents CRLB computation. For the proposed CKM-based methods, the CRLB is computed as the inverse of the FIM, and the FIM is calculated based on the first equation in (15), where the required gradient of the log-likelihood $\nabla_{\mathbf{x}}(-\ln P(\hat{\mathbf{z}}|\mathbf{x}))$ is provided in (41) and is computed efficiently via automatic differentiation. Geometry-based LoS, CKM-based LoS, NLoS, and NLoS_conv are evaluated, with CRLB differences due to paths usage for sensing and distinct sensing functions mapping the relationship between target location and channel knowledge.
@@
The simulation results demonstrate the potential of the proposed CKM-based sensing framework in ISAC systems. By leveraging multi-path information and accurately predicting location-specific channel knowledge using an FCNN, the CKM-based method achieves significant improvements in sensing accuracy in NLoS scenarios compared to traditional geometry-based approaches. The CRLB analysis further confirms the robustness of the CKM-based method to angle and delay estimation errors. These findings highlight the effectiveness of the CKM framework for environment-aware sensing in complex wireless environments.
@@ VI Conclusion
This paper introduces a CKM-enabled ISAC framework that achieves robust environment-aware sensing, even in challenging NLoS scenarios where traditional geometry-based methods fail. We demonstrate that the same CKM designed for environment-aware communications can be exploited to achieve sensing as well. Simulation results demonstrate significant performance gains over geometry-based approaches, with CRLB analysis confirming enhanced robustness. This approach marks a significant advancement for ISAC systems, enabling seamless integration of sensing and communication in complex environments enabled by the same CKM.

### 09 Apr 2026 v4

@@ III Environment-aware ISAC via CKM
where $\mathcal{M}$ represents the CKM, mapping a two-dimensional UE location vector to a channel vector of dimension $D$. CKM exploits these historical pairs $(\mathbf{Q},\mathbf{W})$ to infer new channel knowledge without requiring an explicit characterization of either the mapping function $f(\cdot)$ or the environment $E$. Without loss of generality, CKM captures the statistical relationship between UE locations $\mathbf{q}$ and channel knowledge $\mathbf{w}$, represented as a conditional probability distribution $P_{\mathcal{M}}(\mathbf{w}|\mathbf{q})$. This relationship is learned from the historical data $(\mathbf{Q},\mathbf{W})$ through various methods, such as interpolation, machine learning, or other data-driven techniques, depending on the specific implementation of CKM.
@@
where $h(\cdot)$ is the function taking the composition of functions $f(\cdot)$ and $g(\cdot)$. From (22), the historical data $(\mathbf{Q},\mathbf{W})$ can effectively replace the requirement to explicitly reconstruct the environment $E$.
CKM is obtained by learning from the historical data $(\mathbf{Q},\mathbf{W})$, represented by the mapping $\mathcal{M}$ as shown in (19). Different from explicitly reconstructing the physical environment $E$ from the historical data $(\mathbf{Q},\mathbf{W})$, which is usually computationally prohibitive and unnecessary, CKM serves as an implicit, data-driven representation of the environment $E$. It directly learns the functional “location-to-channel” mapping from historical data, effectively capturing the necessary environment prior knowledge without the overhead of an explicit physical map reconstruction. Therefore, the observation model (22) can be further written as
@@ IV-A Relationship between sensing response and communication channel
A target that reflects signals in all directions is analogous to a virtual communication UE at the same location. This analogy arises from the physical similarity in signal interaction: the target receives the transmitted signal from the BS, attenuates it with a reflection coefficient $\beta$, and reflects it back to the BS. This process mirrors the behavior of a virtual UE that receives the downlink signal and retransmits it via an uplink channel, as illustrated in Fig. 4.
@@ IV-B CADM-enabled sensing
To model the dependence of these parameters on $\mathbf{q}$, CADM employs an FCNN to learn the mapping from the UE location $\mathbf{q}$ to the mean and variance parameters in the Gaussian model in (36). For example, if a two-layer FCNN is used with the input being the UE location $\mathbf{q}\in\mathbb{R}^{2}$, the forward process is defined as
@@
Then, the log-likelihood of the sensing channel knowledge can be written as (40) shown on the top of the next page, where the second equation is due to (32). In (38) and (39), we assume that for a given path $l^{\prime}$, the angle $\theta_{l^{\prime}}$ and delay $\tau_{l^{\prime}}$ are statistically independent. This assumption is based on the underlying physics of multipath propagation in complex environments. From a physical standpoint, the parameters are determined by different geometric properties: The AoA is primarily dictated by the geometric orientation of the final scatterer relative to the receiver array. The path delay, in contrast, is a function of the entire path length traveled by the signal, which may involve multiple different scattering objects before reaching that final scatterer. In the NLoS multipath environment, the factors contributing to the total path length (and thus delay) are numerous and distinct from the geometric angle of the final hop. Therefore, in complex environments, the statistical relationship between the total path length and the final arrival angle is generally weak. It is also noted that the CKM framework is not fundamentally restricted by this assumption. The model could be extended to a full multivariate Gaussian distribution.
@@
where $\mathbf{a}\in\mathbb{R}^{4L^{\prime}\times 1}$ is a selection vector with a 11 at the position corresponding to $\mu_{\theta,l_{T}}(\mathbf{x})$ and 0 elsewhere, $\mathbf{W}_{1}$, $\mathbf{W}_{2}$, $\mathbf{b}_{1}$ are FCNN weights and biases, and $\sigma^{\prime}(\cdot)$ is the derivative of the activation function (such as ReLU). These closed-form expressions enable computation of $\nabla_{\mathbf{x}}(-\ln P(\hat{\mathbf{z}}|\mathbf{x}))$, facilitating optimization of $\mathbf{x}$ via (12). Alternatively, these gradients can be computed efficiently using automatic differentiation in PyTorch by enabling gradient tracking for the location input, calling $backward()$ on the output, and then accessing the input’s .grad.grad attribute. Besides, the $torch.autograd.grad()$ function provides a more direct alternative for this computation.
In summary, to estimate the location $\mathbf{x}$ of the sensing target in the challenging NLoS environment, parameters such as AoD, AoA, and delay for each path are first estimated based on the echo signal $\mathbf{y}(t)$ in (6) using well-established algorithms (e.g., MUSIC, ESPRIT, or compressed sensing [3]), forming the observation vector $\hat{\mathbf{z}}$. The methods cited, such as MUSIC or ESPRIT, are mentioned only as well-established, high-resolution examples, not as mandatory components. For dynamic scenarios involving high-speed targets, where low latency and real-time capability are required, our framework can be seamlessly integrated with various estimation algorithms (e.g., FFT-based) specifically designed for high-mobility tracking. The core contribution of this work lies in intelligently utilizing the outputs $\hat{\mathbf{z}}$ of any such algorithms for NLoS localization. Then, substituting $\hat{\mathbf{z}}$ into (41) allows computation of the gradient, enabling estimation of the target location $\mathbf{x}$ via gradient descent in (12). Note that while the algorithm is presented for the most challenging NLoS scenarios, it is applicable to the LoS case as well. The pseudo-code of CKM-enabled environment-aware sensing is elaborated in Algorithm 1.
@@ V Simulation Results
The simulation is conducted in a 100 m $\times$ 100 m square area, where a BS is located at $(50,0)m$. Besides, there are several scatterers randomly distributed within the area, and one target is placed at a random location for each sensing process. To model the statistical nature of a dynamic environment, we incorporated small random location variations for the scatterers during the data generation process. This ensures that the training data reflects statistical fluctuations, allowing the FCNN to learn a meaningful location-dependent mean and variance for the channel parameters, rather than just a fixed value. The locations of the BS, scatterers, and target are illustrated in Fig. 5, where the BS is marked with a red triangle, scatterers with green circles, and the sensing target with a blue circle. Two sensing scenarios are considered:
@@
To construct the CKM, we use an FCNN as shown in Fig. 6. It consists of five layers: an input layer with 2 neurons (corresponding to the 2D coordinates of UE location $\mathbf{q}$), three hidden layers with $128\times 256\times 128$ neurons using LeakyReLU activation with a negative slope of 0.01, and an output layer with $4L^{\prime}$ neurons, where $L^{\prime}=5$ is the number of Gaussian components in (36). The output includes the angle mean $\mu_{\theta,l^{\prime}}$, the angle variance $\sigma_{\theta,l^{\prime}}^{2}$, the delay mean $\mu_{\tau,l^{\prime}}$, and the delay variance $\sigma_{\tau,l^{\prime}}^{2}$ for each path $l^{\prime}$. Besides, the training dataset is generated by sampling 10,000 points uniformly within the 100 m $\times$ 100 m area. This dataset was empirically found to be sufficient for the FCNN to converge and was partitioned into a training set (80%) and a validation set (20%). Note that in practical deployments, it is unnecessary to physically collect all 10,000 samples. As mentioned in Section I, advanced CKM construction methods based on generative AI [8, 12] can complete a large-scale dataset from only a small fraction of real-world measurements, significantly reducing the data collection burden. For each sampled UE location in the training dataset, the groundtruth channel is generated by considering all possible single-bounce paths via the scatterers. The delay for each path is calculated based on its total geometric length, which is the distance from the BS to the scatterer and then to the UE. The complex gain of each path is determined by two main factors: the path-length-dependent attenuation and a unique random reflection coefficient assigned to each scatterer. Although all paths are generated, the CADM is designed to learn the $L^{\prime}=5$ dominant ones. Therefore, a path selection mechanism is applied: the 5 paths with the highest power, as determined by their path gains, are chosen. The angles and delays of these 5 selected paths constitute the groundtruth data used for training the FCNN, and the training loss is computed as the mean squared error (MSE) between the predicted and true channel parameters. The network was trained for 500 epochs using the Adam optimizer, with an initial learning rate of 0.01 and a batch size of 128.
@@
Fig. 7 presents the RMSE for the location estimation of the sensing target with different methods vs. angle estimation error $\sigma_{\theta}$ and $\sigma_{\phi}$, with fixed delay std $\sigma_{\tau}=20$ ns. It can be seen that the CKM-based methods consistently demonstrate significantly lower RMSE than geometry-based approaches across all angle error levels, by orders of magnitude. Such dramatic performance gains are attributed to the utilization of CADM, which exploits a rich set of NLoS path information, including AoA, AoD and delay. In contrast, even in the mixed LoS/NLoS scenarios, the geometry-based method relies solely on the LoS path’s angle and delay, where angle errors severely degrade performance due to their tight coupling with location as (2). Even with accurate delay estimates, RMSE increases sharply as angle error grows. For the challenging NLoS scenario, the geometry-based method fails to give any meaningful estimation for the sensing target location, since it erroneously assumes the shortest-delay NLoS path as LoS, leading to substantial localization errors. CKM-based methods, by leveraging multiple paths, decouple angle and delay contributions, resulting in much slower RMSE degradation with increasing angle error, as precise delay measurements mitigate the impact of angle inaccuracies. In the pure NLoS scenario, the CKM-based method, employing the proposed generalized sensing model as (6) with up to $L^{\prime 2}$ non-reciprocal paths, outperforms CKM-based NLoS_conv, which uses a conventional model as (5) with only $L^{\prime}$ reciprocal paths. The reduced path diversity in the conventional model limits its information, yielding higher RMSE. Surprisingly, CKM-based NLoS achieves even lower RMSE than CKM-based LoS, especially when the angle errors are large. This may seem counterintuitive at first glance, but is expected because NLoS paths, with distinct AoA, AoD and delay, provide three parameters per path, compared to two for LoS paths (AoA = AoD, delay). The increased path count ($L^{\prime 2}$ vs. $L^{\prime}$) and richer information per path in NLoS scenarios enhance robustness to angle errors when properly exploited through CKM, enabling superior localization accuracy despite the absence of a LoS path. Similar observations can be obtained from Fig. 8, which plots the sensing RMSE as a function of the delay estimation error $\sigma_{\tau}$.
@@
Following the MSE analysis, the theoretical performance bounds are explored through the CRLB to highlight localization performance limits. CRLB is analyzed vs. angle error (with delay error std $\sigma_{\tau}=20$ ns) and delay error $\sigma_{\tau}$ (with angle error std $\sigma_{\theta}=\sigma_{\phi}=0.10$ rad) as shown in Fig. 9 and 10, respectively. The geometry-based NLoS is excluded, as unknown scatterer location prevents CRLB computation. For the proposed CKM-based methods, the CRLB is computed as the inverse of the FIM, and the FIM is calculated based on the first equation in (15), where the required gradient of the log-likelihood $\nabla_{\mathbf{x}}(-\ln P(\hat{\mathbf{z}}|\mathbf{x}))$ is provided in (41) and is computed efficiently via automatic differentiation. The geometry-based LoS, CKM-based LoS, NLoS, and NLoS_conv are evaluated, with CRLB differences due to different path utilization for sensing and distinct sensing functions mapping the relationship between target location and channel knowledge.
@@
The simulation results demonstrate the potential of the proposed CKM-based sensing framework in ISAC systems. By leveraging multipath information and accurately predicting location-specific channel knowledge using an FCNN, the CKM-based method achieves significant improvements in sensing accuracy in NLoS scenarios compared to traditional geometry-based approaches. The CRLB analysis further confirms the robustness of the CKM-based method to angle and delay estimation errors. These findings highlight the effectiveness of the CKM framework for environment-aware sensing in complex wireless environments.
@@ VI Conclusion
This paper introduces a CKM-enabled ISAC framework that achieves robust environment-aware sensing, even in challenging NLoS scenarios where traditional geometry-based methods fail. We demonstrate that the same CKM designed for environment-aware communications can be exploited to achieve sensing as well. Simulation results demonstrate significant performance gains over geometry-based approaches, with CRLB analysis confirming enhanced robustness. This approach marks a significant advancement for ISAC systems, enabling seamless integration of sensing and communication in complex environments using the same CKM.